sally has 210 marbles some are red some are green there are 10 more green marbles how many red ones does she have
Answer by stanbon(57410)

sally has 210 marbles some are red some are green there are 10 more green marbles how many red ones does she have
-------
Equations:
r + g = 210
g = r + 10
----
Substitute for "g" and solve for "r":
r + r+10 = 210
2r = 200
red = 100
===============
cheers,
Stan H.
===========